OpenStudy (anonymous):

Evaluate the expression 3 (y + 1) For y = 2

OpenStudy (anonymous):

A 6 B 9 C 7 D 8

OpenStudy (anonymous):

what do you think you should do first

OpenStudy (er.mohd.amir):

put value of y in equation and solve

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

6

OpenStudy (er.mohd.amir):

no 6 is wrong

OpenStudy (anonymous):

remember the 1 what do you do with it

OpenStudy (anonymous):

add it to 2 to het 3 than 3 +3=6 right

OpenStudy (er.mohd.amir):

3(y+1) =3(2+1) and braket means multiply

OpenStudy (anonymous):

so 2 times 1 = 2 2+3= 5 but there is no 5

OpenStudy (anonymous):

@GabeBae @Loser66

OpenStudy (gabebae):

yes?

OpenStudy (anonymous):

is the answer a 6

OpenStudy (anonymous):

3 x (2 +1) = you do the parentheses first.

OpenStudy (anonymous):

Your first step is to replace y with 2.

OpenStudy (anonymous):

no 6 is incorrect.

OpenStudy (er.mohd.amir):

3(3) 3*3

OpenStudy (anonymous):

so the answer is 9

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

I can help with more questions, if you have them. :)

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

let me just close this one and open a new one ok